# Cell-cell communication inference and analysis: biological mechanisms, computational approaches, and future opportunities

### v1

@@ Abstract
In multicellular organisms, cells coordinate their activities through cell-cell communication (CCC), which are crucial for development, tissue homeostasis, and disease progression. Recent advances in single-cell and spatial omics technologies provide unprecedented opportunities to systematically infer and analyze CCC from these omics data, either by integrating prior knowledge of ligand-receptor interactions (LRIs) or through de novo approaches. A variety of computational methods have been developed, focusing on methodological innovations, accurate modeling of complex signaling mechanisms, and investigation of broader biological questions. These advances have greatly enhanced our ability to analyze CCC and generate biological hypotheses. Here, we introduce the biological mechanisms and modeling strategies of CCC, and provide a focused overview of more than 140 computational methods for inferring CCC from single-cell and spatial transcriptomic data, emphasizing the diversity in methodological frameworks and biological questions. Finally, we discuss the current challenges and future opportunities in this rapidly evolving field.
@@ 1 Introduction
Cell-cell communication (CCC) refers to the process by which cells exchange signals to coordinate their activities through soluble signaling molecules or direct physical contact. It forms the fundamental basis for maintaining homeostasis and achieving functional integration in multicellular organisms. This process involves a series of highly specific molecular mechanisms, requiring the coordinated participation of various molecules—such as ligands, receptors, ions, and metabolites—as well as membrane structures, including tight junctions, gap junctions, and immunological synapses[atakan2014molecular, heldin2016signals, alberts2022molecular, LimMayer2024, clark2018biology]. Elucidating the mechanisms underlying CCC not only deepens our understanding of development and disease, but also identifies potential targets for therapeutic interventions[CCCreviewFan2020, CCCreviewSu2024].
@@
Traditionally, the study of CCC has relied heavily on labor-intensive experimental approaches, such as histological tissue section analysis, in vitro co-culture systems, and in vivo genetic manipulations. These methods are often time-consuming and technically demanding, typically permitting the investigation of communication between only a limited number of cell types and signals, thereby hindering a comprehensive elucidation of CCC networks. The rapid advancement of single-cell RNA sequencing (scRNA-seq) technologies has revolutionized CCC research[CCCreviewFan2020, RN448]. The scRNA-seq data capture gene expression patterns relevant to CCC, providing indirect insights into the abundance of proteins involved in signal transduction. Building on this foundation, numerous computational tools have been developed to infer CCC by integrating scRNA-seq data with prior knowledge of ligand-receptor interactions (LRIs)[CCCreviewFan2020, RN448, RN449, CCCreviewJin2022, CCCreviewNie2023, RN447, CCCreviewSu2024, cesaro2025advances], thus yielding biologically interpretable insights. However, spatial information on cells is inherently lost in scRNA-seq data. Given that most CCC events occur within a short spatial distance, the emerging spatial transcriptomics bring new opportunities to improve CCC analysis[STreview2021integrating, STyuan2021]. The ongoing technological progress, such as single-cell and spatial multi-omics[spatialomicsScience2023, reviewMultiOmics2023], and increasingly complex biological questions continue to drive the evolution of computational approaches, enabling more comprehensive and in-depth analyses.
@@
This review aims to equip biologically focused researchers with practical guidance on when and how to apply existing CCC tools, while providing computational biologists with the biological foundations and methodological advances of CCC to facilitate the development of novel CCC analysis tools. While related reviews exist[CCCreviewFan2020, RN448, RN449, CCCreviewJin2022, CCCreviewNie2023, RN447, CCCreviewSu2024, cesaro2025advances], here we emphasis the biological mechanisms and modeling strategies that underpin CCC, and critically examine the most recent and innovative computational approaches. This review begins with a systematic overview of the biological foundations of CCC, including various types of intercellular and intracellular signaling mechanisms. We then present a general modeling framework for CCC inference, and review a range of computational methods developed in recent years, with an emphasis on their methodological diversification and the specific biological questions they aim to address. Finally, we discuss the key challenges and emerging opportunities in this rapidly advancing field.
@@ 2.1 Types of Cell-Cell Communication
Precise, reliable, and reproducible mechanisms of CCC have evolved to coordinate biological activities among cells within multicellular organisms. Distinct communication systems operate across various tissues and organs to transmit information from signaling cells to their target cells[alberts2022molecular, LimMayer2024, atakan2014molecular]. The chemical signals used by signaling cells for information transmission are known as ligands, while the corresponding binding molecules on target cells are called receptors. Ligands comprise diverse molecules—including proteins (e.g., growth factors, cytokines), peptides, small molecules (e.g., metabolites, neurotransmitters, hormones), and lipids (e.g., prostaglandins, eicosanoids)—that bind to cell-surface receptors. A separate class of hydrophobic ligands (e.g., steroids, gases) diffuses across the plasma membrane to bind intracellular receptors. Receptors are broadly classified by localization and function: cell-surface (transmembrane) receptors and intracellular (cytoplasmic/nuclear) receptors. Cell-surface receptors are further categorized as ion channel-linked, G-protein-coupled, or enzyme-linked[heldin2016signals]. Upon ligand binding, receptors undergo conformational changes that activate intracellular signaling molecules within target cells, and further initiate signal transduction cascades.
CCC can be broadly classified into five types based on underlying principles: paracrine, autocrine, contact-dependent, synaptic, and endocrine signaling (Figure 1). These categories are primarily distinguished by the distance over which ligands act to reach their target cells[clark2018biology, LimMayer2024, atakan2014molecular].
Figure 1: Diverse types of cell-cell communication. Paracrine signaling involves signaling molecules that act on neighboring cells within a localized area by diffusing through the extracellular space. Autocrine signaling refers to signals that act on the same cell that produces them or on a population of identical cell types. Contact-dependent signaling relies on direct physical contact between neighboring cells, such as direct membrane contact or gap junctions. Synaptic signaling enables long-distance communication through the specialized synaptic structures of neurons. Endocrine signaling involves signaling molecules produced by signaling cells that reach distant target cells via extracellular fluids, such as blood. Additionally, extracellular vesicles (EVs) also mediate intercellular signaling by delivering a range of bioactive molecules (e.g., proteins, lipids, nucleic acids) through paracrine, autocrine, and endocrine mechanisms. This figure is inspired by a previous work[alberts2022molecular]. Created in https://BioRender.com
@@
Contact-dependent signaling, also known as juxtacrine signaling, occurs through direct physical contact between the signaling and target cells[fagotto1996cell] (Figure 1). This form of communication can be classified into three major types. The first involves membrane-bound signaling molecules on the surface of the signaling cell that specifically bind to receptors on the plasma membrane of an adjacent target cell. The second type entails the direct delivery of ligands from the signaling cell into the intracellular space of the target cell, where they subsequently interact with receptors. The third mechanism involves the formation of intercellular channels that facilitate the direct transfer of small signaling molecules between the cytoplasms of neighboring cells. Well-characterized examples include gap junctions in animal cells and plasmodesmata in plant cells. Notably, large macromolecules such as proteins and nucleic acids are generally unable to pass through these channels.
@@
The signal transduction mechanisms described above primarily involve short-range communication. However, long-distance CCC is essential for coordinating organism-wide biological activities, as illustrated in Figure 1. The human nervous system represents the most sophisticated form of long-distance signaling, comprising vast networks of neurons that communicate via specialized synaptic structures. Synaptic signaling enables rapid, long-distance, yet contact-dependent communication between neurons[kornberg2014communicating]. Although the cell bodies of communicating neurons may be distant, their elongated axons and dendrites transport signals directly to specialized sites of cell-cell contact. This architecture ensures efficient, spatiotemporally precise information transfer regardless of the distances between the cell bodies of contacting cells. A synapse is a highly specialized structure for information transfer, comprising a presynaptic membrane, a synaptic cleft, and a postsynaptic membrane (Figure 1). Upon arrival of an electrical impulse at the synaptic terminal, neurotransmitters are released from synaptic vesicles into the cleft. These diffusing neurotransmitters then bind to receptors on the postsynaptic membrane, triggering a new electrical signal in the target cell. This process mediates a critical bidirectional conversion between electrical and chemical signaling.
@@
Another major form of long-distance signaling is endocrine signaling (Figure 1). In this process, endocrine cells secrete hormones that are transported through the bloodstream to distant target cells to convey regulatory information. Unlike synaptic signaling, which transmits biological information within milliseconds via direct neuron-to-neuron communication, endocrine signaling depends on circulatory distribution, resulting in significantly slower transmission speeds[desjardins1981endocrine]. The signaling molecules involved in this pathway—collectively known as hormones—are typically characterized by low water solubility, a property that facilitates their long-distance transport in the blood. In contrast, water-soluble signaling molecules are generally involved in paracrine signaling. Despite its slower onset, endocrine signaling exerts more prolonged effects on target cells. Even a small quantity of hormone molecules can elicit sustained regulatory responses, playing a vital role in the modulation of diverse physiological processes.
@@
Notably, beyond classical ligand-receptor interactions, extracellular vesicle (EV)-mediated CCC has garnered significant interest for its potential applications as novel disease biomarkers and drug delivery systems. EVs are membrane-bound vesicles released by cells via plasma membrane budding and fission or multivesicular body formation, encapsulating diverse bioactive molecules such as proteins, lipids, and nucleic acids. Once released into the extracellular space, EVs interact with the extracellular and pericellular matrix, facilitating their transport to recipient cells through autocrine, paracrine, or endocrine routes[van2022challenges]. Upon reaching a target cell, EVs exert their influence through three primary mechanisms: direct binding to cell-surface receptors, fusion with plasma membrane to deliver cargo, or internalization into the cytoplasm [gurung2021exosome]. Unlike conventional signaling, EV-mediated communication enables the coordinated delivery of multiple signaling molecules, potentially eliciting more complex, pleiotropic responses in recipient cells[liu2023review]. This multifaceted interaction repertoire introduces an additional layer of complexity to CCC networks.
@@ 2.2 Signal Transduction in Cell-Cell Communication
Figure 2: Cellular signal transduction mediated by different receptor types and major modes of intracellular signaling crosstalk. (a) Cell-surface receptor signaling: Ligand binding activates a cascade of cytoplasmic relay proteins and transcription factors, resulting in the nuclear translocation of transcription factors to regulate target gene expression. (b) Cytoplasmic receptor signaling: The ligand diffuses across the plasma membrane and binds its receptor in the cytoplasm. The resulting ligand-receptor complex then translocates to the nucleus, where it binds specific DNA regulatory sequences to direct transcriptional regulation. (c) Nuclear receptor signaling: The ligand diffuses across the plasma membrane and into the nucleus, where it binds a nuclear receptor. The ligand-receptor complex then directly modulates gene transcription. (d) Single-pathway activation: A lone intracellular signaling pathway is initiated by ligand-receptor binding. (e) Multiple-pathway activation: A single ligand-receptor pair triggers several distinct signaling pathways. (f) Signal convergence: Multiple, distinct ligand-receptor pairs activate a common downstream signaling pathway. (g) Pathway crosstalk: Interdependent signaling pathways, triggered by different ligand-receptor pairs, mutually influence one another. This figure is inspired by a previous work[Avior2013] and Pearson Education, Inc. Created in https://BioRender.com
@@
The signaling cascade involves intracellular signal propagation. Rather than proceeding as a simple linear chain, intracellular signaling operates as a complex regulatory network composed of various signaling molecules. These molecules can be broadly classified into second messengers and signaling proteins. Second messengers are small intracellular molecules that mediate rapid and broad signal dissemination. In contrast, signaling proteins are large macromolecules primarily responsible for highly specific signal transduction events. Upon receptor activation, second messengers transmit signals by rapidly altering their intracellular concentrations, thereby modulating the activity of downstream effectors. Signaling proteins, due to their large molecular size, undergo minimal concentration changes but exert their effects by activating downstream signaling components or modulating second messenger levels[newton2016second].
@@
Furthermore, during intracellular signal transduction, the addition or removal of phosphate groups plays a critical role in regulating most signaling processes. As a result, signaling proteins often function as molecular switches. Two major classes of these switches are proteins modified through phosphorylation of specific amino acid residues by upstream kinases, and G proteins, which function by cycling between an active GTP-bound state and an inactive GDP-bound state. In addition, specific interactions among signaling proteins further facilitate the transmission and modulation of intracellular signals. Ultimately, the signaling cascade regulates the expression of target genes via activated TFs, thereby inducing a range of biological responses, including cell proliferation, differentiation, and metabolic reprogramming[nishi2015crosstalk].
@@ 2.3 Signal Crosstalk in Cell-Cell Communication
Signaling pathways do not operate in isolation but function within an integrated network of protein–protein interactions. The signaling mode in which intracellular pathways influence one another is referred to as signaling crosstalk, which helps to integrate signals from multiple inputs in different ways, giving rise to the vast range of cellular responses. Signaling crosstalk typically manifests in several distinct patterns[RN442, RN444, RN445] (Figure 2d-g). The first pattern is single signal response, wherein ligand-receptor binding activates only one intracellular signaling pathway (Figure 2d). The second pattern is signal bifurcation, in which a single ligand-receptor pair simultaneously activates multiple signaling pathways (Figure 2e). For instance, fibroblast growth factors (FGFs) can activate the RTK pathway, the STAT pathway, and an additional pathway involving lipid metabolism remodeling and increased intracellular calcium levels[RN439]. The third pattern is convergent signaling, where distinct ligand-receptor pairs converge to activate a shared intracellular pathway, such as during T lymphocyte differentiation[RN439] (Figure 2f). The fourth pattern is modulatory crosstalk, where separate pathways initiated by different ligand-receptor pairs interact such that one pathway modulates—either enhances or inhibits—the transmission of another (Figure 2g). For example, Hh signaling can potentiate Wnt pathway activity, while Wnt signaling, in turn, modulates Hh effectors—a dynamic interplay essential in tissue regeneration and cancer progression[orzechowska2023sonic].

### v3

@@ Abstract
In multicellular organisms, cells coordinate their activities through cell-cell communication (CCC), which is crucial for development, tissue homeostasis, and disease progression. Recent advances in single-cell and spatial omics technologies provide unprecedented opportunities to systematically infer and analyze CCC from these omics data, either by integrating prior knowledge of ligand-receptor interactions (LRIs) or through de novo approaches. A variety of computational methods have been developed, focusing on methodological innovations, accurate modeling of complex signaling mechanisms, and investigation of broader biological questions. These advances have greatly enhanced our ability to analyze CCC and generate biological hypotheses. Here, we introduce the biological mechanisms and modeling strategies of CCC, and provide a focused overview of more than 140 computational methods for inferring CCC from single-cell and spatial transcriptomic data, emphasizing the diversity in methodological frameworks and biological questions. Finally, we discuss the current challenges and future opportunities in this rapidly evolving field, and summarize available methods in an interactive online resource (https://cellchat.whu.edu.cn) to facilitate more efficient method comparison and selection.
@@ 1 Introduction
Cell-cell communication (CCC) refers to the process by which cells exchange signals to coordinate their activities through soluble signaling molecules or direct physical contact. It forms the fundamental basis for maintaining homeostasis and achieving functional integration in multicellular organisms. This process involves a series of highly specific molecular mechanisms, requiring the coordinated participation of various molecules—such as ligands, receptors, ions, and metabolites—as well as membrane structures, including tight junctions, gap junctions, and immunological synapses[15, 73, 3, 109, 39]. Elucidating the mechanisms underlying CCC not only deepens our understanding of development and disease, but also identifies potential targets for therapeutic interventions[152, 160].
@@
Traditionally, the study of CCC has relied heavily on labor-intensive experimental approaches, such as histological tissue section analysis, in vitro co-culture systems, and in vivo genetic manipulations. These methods are often time-consuming and technically demanding, typically permitting the investigation of communication between only a limited number of cell types and signals, thereby hindering a comprehensive elucidation of CCC networks. The rapid advancement of single-cell RNA sequencing (scRNA-seq) technologies has revolutionized CCC research[152, 13]. The scRNA-seq data capture gene expression patterns relevant to CCC, providing indirect insights into the abundance of proteins involved in signal transduction. Building on this foundation, numerous computational tools have been developed to infer CCC by integrating scRNA-seq data with prior knowledge of ligand-receptor interactions (LRIs)[152, 13, 5, 87, 179, 9, 160, 31], thus yielding biologically interpretable insights. However, spatial information on cells is inherently lost in scRNA-seq data. Given that most CCC events occur within a short spatial distance, the emerging spatial transcriptomics brings new opportunities to improve CCC analysis[117, 49]. The ongoing technological progress, such as single-cell and spatial multi-omics[22, 171], and increasingly complex biological questions continue to drive the evolution of computational approaches, enabling more comprehensive and in-depth analyses.
@@
This review aims to equip biologically focused researchers with practical guidance on when and how to apply existing CCC tools, while providing computational biologists with the biological foundations and methodological advances of CCC to facilitate the development of novel CCC analysis tools. While related reviews exist[152, 13, 5, 87, 179, 9, 160, 31], here we emphasize the biological mechanisms and modeling strategies that underpin CCC, and critically examine the most recent and innovative computational approaches. This review begins with a systematic overview of the biological foundations of CCC, including various types of intercellular and intracellular signaling mechanisms. We then present a general modeling framework for CCC inference, and review a range of computational methods developed in recent years, with an emphasis on their methodological diversification and the specific biological questions they aim to address. Finally, we discuss key challenges and emerging opportunities in this rapidly advancing field. To aid the research community, we provide an interactive, web-based summary of all reviewed computational methods (available at https://cellchat.whu.edu.cn). This resource details each method’s overview, computational principles, and addressed biological questions to facilitate efficient method comparison and selection. It features interactive filtering based on methodological characteristics and includes a submission portal for researchers to contribute new tools, ensuring the resource remains current.
@@ 2.1 Types of Cell-Cell Communication
Precise, reliable, and reproducible mechanisms of CCC have evolved to coordinate biological activities among cells within multicellular organisms. Distinct communication systems operate across various tissues and organs to transmit information from signaling cells to their target cells[3, 109, 15]. The chemical signals used by signaling cells for information transmission are known as ligands, while the corresponding binding molecules on target cells are called receptors. Ligands comprise diverse molecules—including proteins (e.g., growth factors, cytokines), peptides, small molecules (e.g., metabolites, neurotransmitters, hormones), and lipids (e.g., prostaglandins, eicosanoids)—that bind to cell-surface receptors. A separate class of hydrophobic ligands (e.g., steroids, gases) diffuses across the plasma membrane to bind intracellular receptors. Receptors are broadly classified by localization and function: cell-surface (transmembrane) receptors and intracellular (cytoplasmic or nuclear) receptors. Cell-surface receptors are further categorized as ion channel-linked, G-protein-coupled, or enzyme-linked[73]. Upon ligand binding, receptors undergo conformational changes that activate intracellular signaling molecules within target cells, and further initiate signal transduction cascades.
CCC can be broadly classified into five types based on underlying principles: paracrine, autocrine, contact-dependent, synaptic, and endocrine signaling (Figure 1). These categories are primarily distinguished by the distance over which ligands act to reach their target cells[39, 109, 15].
Figure 1: Diverse types of cell-cell communication. Paracrine signaling involves signaling molecules that act on neighboring cells within a localized area by diffusing through the extracellular space. Autocrine signaling refers to signals that act on the same cell that produces them or on a population of identical cell types. Contact-dependent signaling relies on direct physical contact between neighboring cells, such as direct membrane contact or gap junctions. Synaptic signaling enables long-distance communication through the specialized synaptic structures of neurons. Endocrine signaling involves signaling molecules produced by signaling cells that reach distant target cells via extracellular fluids, such as blood. Additionally, extracellular vesicles (EVs) also mediate intercellular signaling by delivering a range of bioactive molecules (e.g., proteins, lipids, nucleic acids) through paracrine, autocrine, and endocrine mechanisms. This figure is inspired by a previous work[3]. Created in BioRender. S. Jin, 2026, https://BioRender.com/trsm0kl.
@@
Contact-dependent signaling, also known as juxtacrine signaling, occurs through direct physical contact between the signaling and target cells[52] (Figure 1). This form of communication can be classified into three major types. The first involves membrane-bound signaling molecules on the surface of the signaling cell that specifically bind to receptors on the plasma membrane of an adjacent target cell. The second type entails the direct delivery of ligands from the signaling cell into the intracellular space of the target cell, where they subsequently interact with receptors. The third mechanism involves the formation of intercellular channels that facilitate the direct transfer of small signaling molecules between the cytoplasms of neighboring cells. Well-characterized examples include gap junctions in animal cells and plasmodesmata in plant cells. Notably, large macromolecules such as proteins and nucleic acids are generally unable to pass through these channels.
@@
The signal transduction mechanisms described above primarily involve short-range communication. However, long-distance CCC is essential for coordinating organism-wide biological activities, as illustrated in Figure 1. The human nervous system represents the most sophisticated form of long-distance signaling, comprising vast networks of neurons that communicate via specialized synaptic structures. Synaptic signaling enables rapid, long-distance, yet contact-dependent communication between neurons[98]. Although the cell bodies of communicating neurons may be distant, their elongated axons and dendrites transport signals directly to specialized sites of cell-cell contact. This architecture ensures efficient, spatiotemporally precise information transfer regardless of the distances between the cell bodies of contacting cells. A synapse is a highly specialized structure for information transfer, comprising a presynaptic membrane, a synaptic cleft, and a postsynaptic membrane (Figure 1). Upon arrival of an electrical impulse at the synaptic terminal, neurotransmitters are released from synaptic vesicles into the cleft. These diffusing neurotransmitters then bind to receptors on the postsynaptic membrane, triggering a new electrical signal in the target cell. This process mediates a critical bidirectional conversion between electrical and chemical signaling.
@@
Another major form of long-distance signaling is endocrine signaling (Figure 1). In this process, endocrine cells secrete hormones that are transported through the bloodstream to distant target cells to convey regulatory information. Unlike synaptic signaling, which transmits biological information within milliseconds via direct neuron-to-neuron communication, endocrine signaling depends on circulatory distribution, resulting in significantly slower transmission speeds[45]. The signaling molecules involved in this pathway—collectively known as hormones—are typically characterized by low water solubility, a property that facilitates their long-distance transport in the blood. In contrast, water-soluble signaling molecules are generally involved in paracrine signaling. Despite its slower onset, endocrine signaling exerts more prolonged effects on target cells. Even a small quantity of hormone molecules can elicit sustained regulatory responses, playing a vital role in the modulation of diverse physiological processes.
@@
Notably, beyond classical ligand-receptor interactions, extracellular vesicle (EV)-mediated CCC has garnered significant interest for its potential applications as novel disease biomarkers and drug delivery systems. EVs are membrane-bound vesicles released by cells via plasma membrane budding and fission or multivesicular body formation, encapsulating diverse bioactive molecules such as proteins, lipids, and nucleic acids. Once released into the extracellular space, EVs interact with the extracellular and pericellular matrix, facilitating their transport to recipient cells through autocrine, paracrine, or endocrine routes[169]. Upon reaching a target cell, EVs exert their influence through three primary mechanisms: direct binding to cell-surface receptors, fusion with plasma membrane to deliver cargo, or internalization into the cytoplasm [66]. Unlike conventional signaling, EV-mediated communication enables the coordinated delivery of multiple signaling molecules, potentially eliciting more complex, pleiotropic responses in recipient cells[114]. This multifaceted interaction repertoire introduces an additional layer of complexity to CCC networks.
@@ 2.2 Signal Transduction in Cell-Cell Communication
Figure 2: Cellular signal transduction mediated by different receptor types and major modes of intracellular signaling crosstalk. (a) Cell-surface receptor signaling: Ligand binding activates a cascade of cytoplasmic relay proteins and transcription factors, resulting in the nuclear translocation of transcription factors to regulate target gene expression. (b) Cytoplasmic receptor signaling: The ligand diffuses across the plasma membrane and binds its receptor in the cytoplasm. The resulting ligand-receptor complex then translocates to the nucleus, where it binds specific DNA regulatory sequences to direct transcriptional regulation. (c) Nuclear receptor signaling: The ligand diffuses across the plasma membrane and into the nucleus, where it binds a nuclear receptor. The ligand-receptor complex then directly modulates gene transcription. (d) Single-pathway activation: A lone intracellular signaling pathway is initiated by ligand-receptor binding. (e) Multiple-pathway activation: A single ligand-receptor pair triggers several distinct signaling pathways. (f) Signal convergence: Multiple, distinct ligand-receptor pairs activate a common downstream signaling pathway. (g) Pathway crosstalk: Interdependent signaling pathways, triggered by different ligand-receptor pairs, mutually influence one another. This figure is inspired by a previous work[16] and Pearson Education, Inc. Created in BioRender. S. Jin, 2026, https://BioRender.com/xa3ut41.
@@
The signaling cascade involves intracellular signal propagation. Rather than proceeding as a simple linear chain, intracellular signaling operates as a complex regulatory network composed of various signaling molecules. These molecules can be broadly classified into second messengers and signaling proteins. Second messengers are small intracellular molecules that mediate rapid and broad signal dissemination. In contrast, signaling proteins are large macromolecules primarily responsible for highly specific signal transduction events. Upon receptor activation, second messengers transmit signals by rapidly altering their intracellular concentrations, thereby modulating the activity of downstream effectors. Signaling proteins, due to their large molecular size, undergo minimal concentration changes but exert their effects by activating downstream signaling components or modulating second messenger levels[131].
@@
Furthermore, during intracellular signal transduction, the addition or removal of phosphate groups plays a critical role in regulating most signaling processes. As a result, signaling proteins often function as molecular switches. Two major classes of these switches are proteins modified through phosphorylation of specific amino acid residues by upstream kinases, and G proteins, which function by cycling between an active GTP-bound state and an inactive GDP-bound state. In addition, specific interactions among signaling proteins further facilitate the transmission and modulation of intracellular signals. Ultimately, the signaling cascade regulates the expression of target genes via activated TFs, thereby inducing a range of biological responses, including cell proliferation, differentiation, and metabolic reprogramming[132].
@@ 2.3 Signal Crosstalk in Cell-Cell Communication
Signaling pathways do not operate in isolation but function within an integrated network of protein–protein interactions. The signaling mode in which intracellular pathways influence one another is referred to as signaling crosstalk, which helps to integrate signals from multiple inputs in different ways, giving rise to the vast range of cellular responses. Signaling crosstalk typically manifests in several distinct patterns[150, 123, 74] (Figure 2d-g). The first pattern is single signal response, wherein ligand-receptor binding activates only one intracellular signaling pathway (Figure 2d). The second pattern is signal bifurcation, in which a single ligand-receptor pair simultaneously activates multiple signaling pathways (Figure 2e). For instance, fibroblast growth factors (FGFs) can activate the RTK pathway, the STAT pathway, and an additional pathway involving lipid metabolism remodeling and increased intracellular calcium levels[64]. The third pattern is convergent signaling, where distinct ligand-receptor pairs converge to activate a shared intracellular pathway, such as during T lymphocyte differentiation[64] (Figure 2f). The fourth pattern is modulatory crosstalk, where separate pathways initiated by different ligand-receptor pairs interact such that one pathway modulates—either enhances or inhibits—the transmission of another (Figure 2g). For example, Hh signaling can potentiate Wnt pathway activity, while Wnt signaling, in turn, modulates Hh effectors—a dynamic interplay essential in tissue regeneration and cancer progression[135].